OpenStudy (anonymous):

Solve: 4(2x - 1) - 3 = 9(x - 4) Select one: a. -29 b. 32 c. 13 d. 29 plz help

OpenStudy (anonymous):

start with multiplying out the parenthesis. What I mean is, for example, 4(2x-1) = (4*(2x) - 1*(2x)) so 4(2x-1) = 8x-4. Now do the other side

OpenStudy (anonymous):

you gonna try MINILUXAILI ??

OpenStudy (anonymous):

:OOOO ok i all ready did it thank ye both

OpenStudy (anonymous):

its 29

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

XD